Ticket Number: HEA-NOC/20040809-1 Ticket Status: UPDATE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20040809 10:42 UTC+1 Problem Start: 20040811 08:00 UTC+1
Ticket Update: 20040816 10:39 UTC+1 Problem End:
Site/line: Boston College, Crumlin VEC, DCU, DIAS, DIT, EPA, HEA,
HETAC, HRB, Hermes, IRCSET, ITnet, Kilester VEC, Liberties
VEC, MLE, NCAD, NCIRL, NLI, NQAI, NUI, NUIM, RCSI, RIA,
TCD, Teagasc, Tipp Inst, UCD, Whitehall VEC, mirrors
HEAnet's main Dublin PoP access router requires a chassis replacement. This
work will take place during an extraordinary maintenance window from
10:00-12:00 on Saturday August 14th.
In preparation, work will take place on Wednesday 11th during the
maintenance window from 08:00-10:00 to ensure resilience for serial links
and backup ATM connections.
The main work will cause an extended outage to the primary ATM,
GigabitEthernet and Ethernet connectivity of affected clients.
Any client with concerns regarding the impact of this maintenance should
contact [log in to unmask] (+353 1 6609040).
20040809 10:26 AH Issue ticket.
20040811 09:46 DW Wednesday work complete; a backup ATM route to Citywest
has been provisioned to remove a dependency on the main access router,
which should provide backup for serial and ATM customers. The substantive
maintenance is still expected on Saturday morning.
20040813 15:36 AH Running order confirmed. In order to prioritise uplink
traffic for clients, ftp.heanet.ie will be unavailable during this
20040814 11:28 SY Work completed at 11:10. Will continue to monitor.
20040816 10:38 RD MRTG graphs now being updated for new ifindex values.
This ticket can be monitored at http://www.hea.net/tickets/20040809-1